The candy shack has 20 lb of mixed white and dark chocolates worth 7.50 per pound. White chocolate alone will sell for 8.00 a pound and dark chocolate sell for 6.00 a pound. How many pounds of each are in the mixture
Answer by stanbon(75887)

The candy shack has 20 lb of mixed white and dark chocolates worth 7.50 per pound. White chocolate alone will sell for 8.00 a pound and dark chocolate sell for 6.00 a pound. How many pounds of each are in the mixture
--------------------
Equations:
w + d = 20 lbs
8w + 6d = 7.5*20 dollars
-----------------------------
Modify for elimination:
8w + 8d = 8*20
8w + 6d = 150
---------------
Subtract and solve for "d":
2d = 10
d = 5 lbs (wt. of dark)
---
Solve for "w":
w = 20-d = 15 lbs (wt of white)
-------------
Cheers,
Stan H.
----------